OpenStudy (gracygirl):

Help please? P = 3r + 2s P - 2s = 3r + 2s - 2s P - 2s = 3r Now what? I'm solving for s.

OpenStudy (will.h):

are those options or the question?

OpenStudy (gracygirl):

The question of what I've worked it out so far, the original is the top, and my teacher said to move the 3r to the other side which I did but I don't know what to do next and it's faster to get help on here because it could be hours before my teacher replies.

OpenStudy (will.h):

to get this clear you wanna solve for p = 3r + 2s solve for s?

OpenStudy (gracygirl):

Yup

OpenStudy (will.h):

fantastic. you made a mistake at the 1st step you should've subtracted 3r from both sides instead of 2s

OpenStudy (gracygirl):

Oh true that >.<

OpenStudy (will.h):

like the following p - 3r = 2s and now divide both sides on 2 \[\frac{ p-3r }{ 2 } = s\]

OpenStudy (gracygirl):

And that's the answer?

OpenStudy (will.h):

hope that helps yes that's the answer we solved for s

OpenStudy (gracygirl):

Thank you, that makes more sense now.

OpenStudy (will.h):

you welcome.